I really need elp with his question: h(x) = -x squared plus 6x and g(x) = 70
h(g(x))=

Found 2 solutions by pradhyumna.kumar@gmail.com, stanbon:
Answer by pradhyumna.kumar@gmail.com(62)

h(x) = -x squared plus 6x and g(x) = 70
h(g(x))= h(70)= - (70)^2 + 6 x 70
= -4900 + 420
= - 4480

Answer by stanbon(75887)

h(x) = -x squared plus 6x and g(x) = 70
h(g(x))= h[70]
= 70^2+ 6*70
=4900+420
=5320
Cheers,
Stan H.